E6.1 (LO 1), AP rooms that are rented at $60 a night. Operating costs are as follows. Service The Soma Inn is trying to determine its break-even point. The inn has 75 Salaries Utilities $10,600 per month Maintenance Maid service Other costs 2,400 per month 1,500 per month 800 per month 8 per room 34 per room Instructions a. Determine the inns break-even point in (1) number of rented rooms per month and (2) dollars. b. If the inn plans on renting an average of 50 rooms per day (assuming a 30-day month), what is (1) the monthly margin of safety in dollars and (2) the margin of safety ratio? E6.2 (LO 1), AP Service In the month of June, Jose Heberts Beauty Salon gave 4,000 haircuts, sham- poos, and hair colorings at an average price of $30 each. During the month, fixed costs were $16,800 and variable costs were 75% of sales. Instructions a. Determine the contribution margin in dollars, per unit and as a ratio. b. Using the contribution margin technique, compute the break-even point in dollars and in units. c. Compute the margin of safety in dollars and as a ratio.
Answer #1

Answer

  • All working forms part of the answer
  • Exercise E 6.1

Working

A

Rent Revenue

$60

variable costs:

Maid service

$8

Other costs

$34

B

Total variable cost

$42

C = A - B

Contribution margin per room

$18

D = (C/A) x 100

CM Ratio

30%

--requirement ‘a’

Salaries

$10,600

Utilities

$2,400

Depreciation

$1,500

Maintenance

$800

A

Total Fixed Cost

$15,300

B

Contribution margin per room

$18

C = A/B

Break Even in no. of rented rooms per month

850

D

CM Ratio

30%

E = A/D

Break Even in dollars

$51,000

--Requirement ‘b’

A = 50 rooms x 30 days

Monthly rooms occupied

1500

B

Rent per room

$60

C = A x B

Monthly rent

$90,000

D

Break Even in dollars

$51,000

E = C - D

Monthky Margin of safety in dollar

$39,000

F = (E/C) x 100

Margin of Safety ratio

43.33%

  • Exercise E 6.2

--Requirement ‘a’

A = 4000 x $30

Service revenue

$120,000

B = A x 75%

Variable Cost

$90,000

C = A - B

Contribution margin

$30,000

D

No. of haircuts

$4,000

E = C/D

Contribution margin per unit

$                     7.50

F = (C/A) x 100

CM ratio

25%

--Requirement ‘b’

A

Total Fixed Cost

$16,800

B

CM ratio

25%

C = A/B

Break Even in dollars

$67,200

D

Contribution margin per unit

$                     7.50

E = A/D

Break Even in units

2240

--Requirement ‘c’

A

Total Revenue

$120,000

B

Break Even in dollars

$67,200

C = A - B

margin of safety in dollar

$52,800

D = (C/A) x 100

Margin of Safety ratio %

44%